Operations Report – April 10th

Crew 192 Operations Report 10Apr2018

SOL 2

Name of person filing report: Victoria Varone

Non-nominal systems: None

Notes on non-nominal systems: N/A

Generator (hours run): 13 hours

Solar — SOC 100%

Diesel – 70%

Propane – 68% (telemetry reported) psi.

Ethanol Free Gasoline (5 Gallon containers for ATV) – N/A

Water (trailer) – N/A
Water (static) – 1020 gallons (two external tanks)
Trailer to Static Pump used – No
Water (loft) – Static to Loft Pump used – Yes
Water Meter: 133303.1

Toilet tank emptied: No. Last emptied April 9.

ATV’s Used: Blue #3, Red #1, and Blue #2.

Oil Added? No

ATV Fuel Used: 25%

# Hours the ATVs were Used today: 4

Notes on ATVs:

Deimos rover used: No
Hours: 0
Beginning charge: N/A
Ending charge: N/A
Currently charging: Yes

Spirit rover used: Yes
Hours: 2
Beginning charge: 100%
Ending charge: 100%
Currently charging: No

Opportunity rover used: No
Hours: 2
Beginning charge: 100%
Ending charge: 100%
Currently charging: Yes

Curiosity rover used: No
Hours: 0
Beginning charge: N/A
Ending charge: N/A
Currently charging: No

HabCar used and why, where? No

General notes and comments: A good day overall, the crew is doing very well.

Summary of internet: 569 MB remaining at 9:30PM.

Summary of suits and radios: Suits all charging, 4 one-piece suits used today. Radios working.

Summary of Hab operations: Changed out air filter, constructed minor and non-permanent communications modifications for radio clips in spacesuits, secured step outside of the side airlock.

Summary of GreenHab operations: See GreenHab report.

Summary of ScienceDome operations: None.

Summary of RAM operations: The third large external water tank on the trailer was rolled into the RAM for cleaning during additional EVAs.

Summary of health and safety issues: None.

Questions, concerns, and requests to Mission Support: The hot water faucet in the hab kitchen appears to be faulty, it was difficult to shut off after use. To avoid the loss of water, we’re not going to use it.